Oxford Materials IT Committee Meeting

HILARY TERM 2006
A meeting of the Department of Materials IT Committee was held on Tuesday 31st January 2005.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of 25th October 2005
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 University IT Users Group – brief verbal report (PJW)
  • 6 IT/AV expenditure - matters for report (PJW)
  • 7 AOB
  • 8 Date and time of next meeting: Tuesday 9th May 2006 at 2pm.

ITEM 2. MATTERS ARISING FROM THE MEETING OF 10th MAY 2005

See minutes below.

ITEM 3: IT/AV MATTERS FOR DISCUSSION - 10th MAY 2005

University ICT strategy and restructuring of central IT support services.

ITEM 4: IT/AV HEAD OF IT REPORT – 10th MAY 2005

Admin Network
Several major staff changes have occurred. (GDWS->CRMG; RMP->APD; RAM->TMcA)
Four new computers were installed (Grovenor, Spruce, Davies, Fellows) and their computers were trickled to the library and digital darkroom. Extra computers were installed for GDWS’s PA and ICS.

Teaching Laboratory
The colour printer was replaced with an HP4700dn. The spare mono printer Kyocera-fs1020D was recycled (replacing the HP4500 for PG and LJFJ). Teaching Laboratory laptops have been lent around the department on numerous occasions when requests do not conflict with Teaching Laboratory requirements.

Lecture theatres / meeting rooms
The CETL bid had the equipment fund slashed thus the proposed video-conference link to Begbroke was not funded. A modified bid to the HEFCE e-learning fund was also unsuccessful. LR7 is now a meeting room, so the computer was removed. 21BRLT computer was upgraded by trickledown from the Admin network. Re-wiring of the 21BR committee room is imminent. The roombooking system was migrated to a different server and the software updated.

Digital Darkroom
Problems with the HP8550 colour printer persist. An engineer visited site but was unable to diagnose the problem. AMcK has a manual workaround, but a long term solution is required. One computer was upgraded by trickledown from the Admin network.

Library
Both open-access computers were upgraded by trickledown from the Admin network.

MML
A new fileserver (Rothery) was configured and all users migrated. The older HP workstations have been decommissioned. The test cluster was expanded to 8PC’s and benchtests carried out, following which a full 32 (48?) node cluster has been specified and is being purchased.

EM
Clive Downing was appointed to replace KS and to run the EM IT lab, but KS remains system administrator for some Unix systems (emgroup1, sgi’s) and liases with Clive over various other systems. A method for unprivileged access to the Department-wide Altiris system for backup of EM PC’s is still being investigated.

Website
The intended review of research group websites last summer still has not happened. The websites hosted on databases.materials were migrated to the new webserver(s).

Publications
Research in Progress 2005 was completed and printed and the website updated.

Network
The new department firewall seems to have been working well and a methodology for daily log-checking and refinement of rules is continuing development of the firewall and network monitoring. Network problems on return at the start of January were mitigated by turning off spanning-tree-protocol on all switches. Switches were upgraded in the EM area and the EM IT lab. The FROntDOor project to install active devices in all buildings under the control of OUCS is due to happen before Easter. The re-wiring in the basement of ETB is complete but the network feeds in ETB still need to be re-routed. Re-wiring of 21BR has started again and is likely to continue until August.

Servers
Materials2(main netware fileserver) remains unchanged.
Oums-nomad (Altiris Deployment Server, antivirus CID, FM Group database).
All websites have been transferred off Databases onto either Www-int or Www-ext.
Databases and Resolve2 are almost ready to be decommissioned.
Oums-wsus (windowsupdates) server has been installed and the old sus(windowsupdates) server will be decommissioned when all users have been transferred.
Test installations using Novell OES Linux and Novell OES Netware are being run on Materials0 and Materials1 currently.

Phones
Faraday people (PSG) lost access to phones for three days when telecoms upgraded the exchange-side of the voice-over-ip (VOIP) phones which resulted in the 5 older phones in the Faraday area needing to be replaced.

Security
Three computers were hacked and blocked by OUCS. As a result of this, the firewall rules were strengthened to disallow incoming windows-networking from Oxford as well as the World. One hacked PC was reinstalled and the other two await reinstallation when theses have been completed. Two users were reprimanded for p2p filesharing. Two of the websem PCs are remain blocked at from previously.

Most computers have now been updated to the Sophos AntiVirus Version 5? Clients (that have better reporting tools) and the sophos CID on databases will be decommissioned soon. For certain cases the Version 4 clients will need to remain active (updating from www-int) and it is supported until Feb 2007. Doorman’s CID for 9x computers will be replaced by a CID on www-int.


MINUTES

A meeting of the Department of Materials IT Committee was held on Tuesday 31st January 2006.
The following people attended: APD, PJW, AMC, KS, JAMS, CJS, Ian Atkin (OUCS)

1. Apologies for Absence and Membership of the Committee

Apologies were received from AHM due to an IDSAF meeting being postponed from the earlier in the morning. The committee welcomed Alana Davies (APD) onto the committee in her capacity as the new Department Administrator. The committee welcomed Chris Salter (CJS) onto the committee replacing Crispin Hetherington as an Academic representative. The committee welcomed Ian Atkin as a representative of OUCS.

2. Matters arising from the minutes

a) Lecture theatres / meeting rooms (Minute 2a)
The University-wide CETL project, for which Adrian Taylor is the department contact, was funded but the equipment fund was slashed thus the proposed video-conference link to Begbroke was not funded. A modified bid to a HEFCE e-learning fund was also unsuccessful.

b) Email and shared calendar server (Minute 2b)
Requirements of an Exchange server have been investigated. Informal comments on Lotus Notes as used in Central Admin were unfavourable.

Action: AMC/PJW

c) IT disaster recovery plans (Minute 2c)
PJW is to circulate draft documentation and to start the process of collecting full details. Item was carried to next meeting.

Action: PJW

d) Gigabit networking (Minute 2d)
iii) The microwave link annual maintenance (£7k/year) has been terminated but the equipment remains “hot” while PJW still looks for a possible buyer for the equipment. The radio license will expire in May and is already marked for cancellation.

The cost and technical requirements for a (low bandwidth) backup network connection as part of the disaster recovery planning was investigated.

It is NOT intended to keep the microwave as disaster back-up due to the high annual maintenance cost. If not maintained/used then it is unlikely to be reliable in the long term as a failover-system.

The current scenario in case of failure of the fibre (e.g. breakage due to a road-digger) is that the Oxford Conference room computer, which is already on a 512k ADSL connection, could be used for urgent email/web usage with access being controlled by Begbroke reception. It is not currently simple to swap computers on the Materials subnet over onto the ADSL network.

Technically it is possible to envisage a system whereby the Materials subnet at Begbroke could be routed across an ADSL/SDSL connection and into a VPN server within Materials, but this is considered technically difficult to implement and potentially unreliable as a failure-over system.

An alternative would be to implement radio signal connection from Begbroke to Oxford, similar to the microwave but using cheaper off-the-shelf equipment on a public frequency. This option could be investigated further in deemed necessary.

Currently is proposed NOT to put in place any provision specifically for Materials. The fibre is under OUCS control and effectively part of the back-bone infrastructure and in the event of a break systems are in place for rapid repair as quickly as possible, although in reality such repairs are likely to be a day or two. (example from August 2005 – Comlab was online again within 5hours!)

iv) Also discussed was the possible effect of upgrading the network to Gigabit upon the Quality of Service for connection sensitive applications such as VOIP phones and videoconferencing. PJW was actioned to investigate further.

In the end the new fibre link to Begbroke did not require another backbone connection and upgrading the main department connection to gigabit was deemed unnecessary given the current traffic patterns. Also since the Begbroke optic cable has been active there have been few complaints about poor quality VOIP phonecalls.

However, investigation did determine that whether the network is 100Mbit or 1000Mbit should not really affect QoS for the VOIP phones. VOIP phone calls degrade in quality when the network traffic at some point between the exchange and the phone is sufficiently high usage that a significant proportion of the VOIP packets are lost. The faster network (1000Mbit) could still be saturated by a large data-transfer, but it would be saturated for less time. Currently there are typically 4-5 switches in Materials between the phone and the OUCS backbone so QoS would need to be turned on on all the switches between the phones and the backbone and this would require replacing the core hardware (3300fx switches). The issue of VOIP QoS should be readdressed in the future when planning to upgrade the 3300fx switches, but currently there are few problems being reported with the VOIP phones.

e) Begbroke Directorate IT Support (Minute 2e)
RMP contacted the Begbroke Directorate to discuss arrangements for sharing IT support provision at Begbroke to cover holidays and illness etc. just before leaving but since the new IT support office at Begbroke, Karen James, has more links with Engineering Science it was decided not to share arrangements for cover with Materials.

f) Website (Minute 2f)
The proposed review of research group websites did not happen during the summer but has been carried forward to the next meeting. Advice and guidance is to be offered by PJW upon compliance with University Website Policies and upon mechanisms for easy updating of content. PJW to email notices to raise interest in web authoring and best practice.

Action: PJW

g) Membership (Minute 8a)
Crispin Hetherington resigned from the ITC committee. Crispin was thanked for his contribution to the Committee. Chris Salter was identified as a possible replacement and was appointed by CRMG.

3 Matters for Discussion

University ICT strategy and restructuring of central IT support services were discussed with input from Ian Adkin as a representative of OUCS. The University is currently developing an ICT strategy in order to create a framework for planned development of ICT. Full details are being posted and updated regularly at http://www.ict.ox.ac.uk. The ICT Strategy document is due to be completed in April. Department IT staff are generally aware of progress but not heavily involved in providing input. Central IT support services have been restructured with the formation of an ICT Team to deliver desktop services and support for all users within Central University Admin, Oxford University Library Services (OULS), and Oxford University Computing Services (OUCS) via an Enhanced Computing Environment (ECE). This is initially unlikely to affect Materials however there is expected to be scope for learning from “best practice”.

4. Head of IT Report

A written report was circulated with the agenda (above).

5. Oxford University IT Users Group Meeting http://www.ict.ox.ac.uk/oxford/groups/ouitug/

a) Skype
The rules banning Skpe (voice-over-IP software) have been relaxed and it is now permissible as long as software has been configured according to the setup instructions to be found at: http://www.oucs.ox.ac.uk/network/voip/skype.xml The department policy however will remain that alternative software is preferable.

b) Freshers computer literacy data
It was reported that 81% of freshers have laptops. An idea of trying to get students to buy standardized laptops before coming to Oxford was tried by Chemistry but it was found that many already had their laptops before even starting to apply to Oxford. In Materials only 1 of this year’s intake admitted to not having their own PC and I did not ask whether people had laptops or desktops.

It was reported that 19% of freshers did not know if their PC’s were running anti-virus software. This emphasises the importance of the Colleges and Departments being very pro-active in ensuring all machines run anti-virus software and are kept fully patched.

c) New Library Management System
A new system (to underpin OLIS) is due to go-live at the end of August. It is expected that some changes will be needed in our library, notably in the self-issue system and the Librarian’s PC.

6. IT/AV Expenditure

PJW apologised for not having prepared a status report on IT expenditure. The potential for selling the old microwave equipement was again mentioned (see Item 2 d iii above).

7 AOB

a) Copyright
APD noted that new copyright regulations need to be assessed and staff notified of changes.

Action: PJW

b) Website
JAMS raised several points concerning the department website. The new projects page needs to be clearer for prospective students to find projects of interest. The current status of representation of research groups on the Department website is patchy at best, and a centralized system for summarizing Research Group information is desirable. PJW to liase with JAMS. Ian Atkin queried why all the IT section of the website is currently restricted to only being visible from within Materials. PJW to assess which content (e.g. policies) should be more widely accessible.

Action: PJW/JAMS

8 Date of Next Meeting

The next meeting will take place on Tuesday 9th May 2006 at 2pm

The meeting closed at 3.50pm